OpenStudy (anonymous):

Solve the system by the substitution method. y=-4x-8 2x-3y=10 Give the solution set. I tried the obvious 2x-3(-4x-8)=10 But it hasn't gotten me anywhere, I've been working on this for a long time now though so i'm probably overlooking something easy. Any help?

OpenStudy (campbell_st):

made a typo so if y = -4x - 8 then substituting 2x -3(-4x - 8) = 10 2x + 12x + 24 = 10 solve for x then substitute into y = -4x - 8 to find y

OpenStudy (campbell_st):

the equation in x you need to solve is 14x + 24 = 10 subtract 24 from both sides 14x = -14 x = -1

OpenStudy (anonymous):

Ah! Totally makes sense, I got to that point before but in my frustration -14 divided by 14 = 0. Who knows what I was thinking. Thank you!!